# Model-Based Clustering with Sequential Outlier Identification using the Distribution of Mahalanobis Distances

### v1

@@ 5.2 TCLUST Simulation Results
ContaminatedMixt fits both contaminated and uncontaminated models but for 19 of the 200 data sets in this simulation study, its best model was uncontaminated. This means that for these data sets, ContaminatedMixt did not classify any data points as outliers.
@@
outlierMBC-minimum 0.96 0.94 4.67 7.66 97.01
outlierMBC-backtrack 0.96 0.92 3.25 11.25 92.00
TCLUST ($\alpha=0.1$) 0.95 0.93 7.08 7.08 100.88
mclust 0.94 0.90 21.93 1.50 120.44
OTRIMLE 0.90 0.87 12.26 11.38 100.88
ContaminatedMixt 0.89 0.78 35.52 13.94 121.58
ContaminatedMixt, mclust, and OTRIMLE failed quite dramatically for some of the data sets, whereas outlierMBC was more reliable. The highest number of false positives identified by ContaminatedMixt was 354, compared to 130 for mclust, 75 for OTRIMLE, 18 for outlierMBC-minimum, and 17 for outlierMBC-backtrack. ContaminatedMixt had more than 50 false positives for 32 data sets, while mclust and OTRIMLE exceeded this number 23 times and 4 times, respectively. The highest number of false negatives for ContaminatedMixt and OTRIMLE was 100, compared to 42 for outlierMBC-backtrack, 36 for outlierMBC-backtrack, and 8 for mclust. ContaminatedMixt and OTRIMLE exceeded 50 false negatives 20 times and 11 times, respectively. This shows that outlierMBC was less likely to produce an extreme number of false positives or false negatives.

### v2

@@ 1 Introduction
Our method is implemented in R as a package called outlierMBC (Doherty et al., 2025), which is published on CRAN, the Comprehensive R Archive Network. Code for reproducing this paper’s analyses is available at github.com/UltanPDoherty/outlierMBCpaper_code.
@@ 5.2 TCLUST Simulation Results
ContaminatedMixt fits both contaminated and uncontaminated models but for 20 of the 200 data sets in this simulation study, its best model according to BIC (Bayesian Information Criterion, Schwarz (1978)) was uncontaminated. This means that for these data sets, ContaminatedMixt did not classify any data points as outliers. Similarly, OTRIMLE did not identify any outliers for 14 of the 200 data sets.
@@
outlierMBC-minimum 0.96 0.94 4.50 8.00 96.50
outlierMBC-backtrack 0.96 0.92 3.14 11.75 91.39
TCLUST ($\alpha=0.1$) 0.95 0.93 6.76 6.76 100
mclust 0.94 0.90 25.32 1.02 124.30
ContaminatedMixt 0.89 0.77 32.99 14.26 118.74
OTRIMLE 0.86 0.83 10.13 16.18 93.96
ContaminatedMixt, mclust, and OTRIMLE failed quite dramatically for some of the data sets, whereas outlierMBC was more reliable. The highest number of false positives identified by ContaminatedMixt was 295, compared to 204 for mclust, 67 for OTRIMLE, 22 for TCLUST, 14 for outlierMBC-minimum, and 13 for outlierMBC-backtrack. ContaminatedMixt and mclust had more than 50 false positives for 37 data sets and 25 data sets, respectively, while OTRIMLE only exceeded this number once. ContaminatedMixt and OTRIMLE exceeded 50 false negatives 22 times and 25 times, respectively, compared to twice for outlierMBC-backtrack and once for outlierMBC-minimum. This shows that outlierMBC was less likely to produce an extreme number of false positives or false negatives.